# A Survey on Human-Centered Evaluation of Explainable AI Methods in Clinical Decision Support Systems

### v3

@@ 3.3 CDSS adopting XAI Methodologies in Deep Learning
Finally, across various clinical specialties, studies consistently highlight the importance of human-centered evaluation in determining the utility of XAI methods. In Ophthalmologysingh2021evaluation assessed multiple visual explanation techniques for a CNN-based model trained to diagnose retinal diseases using a counterfactual explanatory method, showing positive reception by radiologists. While in Cardiologyneves2021interpretable similarly applied various explainability methods to a CDSS based on ECG time series data for predicting heart arrhythmias. In both domains, surveys identified attribution methods, specifically DeepTaylor and LIME, as particularly useful, especially for less experienced clinicians. In Critical Carezhang2024rethinking developed SepsisLab, leveraging an attention module (explanatory part) feeding into an LSTM to highlight important variables for sepsis prediction. Here, semi-structured interviews revealed that while explanations were valued, shifting the AI focus away from final decision could further strengthen the user experience. Next, in Dermatologychanda2024dermatologist implemented GradCAM visualizations for CDSS assisting to diagnose skin melanoma. The system was evaluated with 116 clinicians across multiple phases. Although XAI did not significantly improve diagnostic accuracy, it notably increased clinicians’ confidence and trust. Finally, in Sleep Medicinehwang2022clinical developed a CNN-based CDSS to measure sleep staging, providing explanations through convolutional filters, saliency maps, and intermediate activations. User studies indicated that explanations effectively confirmed AI predictions but were less helpful in determining ambiguous cases, reflecting the clinicians’ desire to validate predictions while relying on their domain expertise.
@@
Of the 31 studies, we identified 14 studies leveraging deep learning for CDSS development. CAM-based methods (including GradCAM) were the most adopted XAI framework with 5 occurrences. However, we documented a scattered application of several frameworks ranging from integrated gradients to LIME, with no clear framework to dominate as compared to SHAP for CDSS evaluated in the previous section. As for clinicians’ perception, the majority proved to helpful with 11 studies showing positive reception ellenrieder2023promoting; pumplun2023bringing; sabol2020explainable; kayadibi2025ai; singh2021evaluation; neves2021interpretable; chanda2024dermatologist; singla2023medical; ihongbe2024evaluating. The remaining papers documented neutral XAI perceptions with:  cabitza2025fromoracular finding that less experienced practitioners found explanations cognitively demanding, zhang2024rethinking shifting XAI away from final decision was perceived more reliable, and  hwang2022clinical finding that AI predictions were less helpful in revealing ambiguous cases, reflecting the clinicians’ preference to validate predictions while relying on their domain expertise.
@@ 4.1 Thematic Analysis of XAI-based CDSS
Next, we briefly outline the main evaluation themes that emerged, reporting a summary in Table 2. Firstly, positive perceptions were mostly linked to Theme 1: Improvements in Risk Assessments. Here, most of the works surveyed fall in this theme, e.g., brennan2019comparing; he2024vms; hur2025comparison; kumarakulasinghe2020evaluating; jing2025development; famiglini2024evidence; kayadibi2025ai, among others, where XAI was seen as an augmentative assisting tool to clinicians to improve their diagnoses. In this theme, we identified sub-themes such as: (1.1) Support for Less Experienced Clinicians neves2021interpretable, where junior clinicians perceived improvements in risk assessment, (1.2) Reduction of False Learning ellenrieder2023promoting, where clinicians were critically able to spot erroneous recommendations, and (1.3) Enhancement of Clinical Planning abraham2023integrating, implying that doctors could better schedule treatments regarding their patients.
@@
However, incorporating XAI may still pose challenges and limitations. Therefore, we discuss themes that emerged from studies expressing neutral perceptions of XAI. Firstly, several works align with Theme 2: Preference for Established Clinical Practice (kovalchuk2022three; matthiesen2021clinician; zhang2024rethinking; sivaraman2023ignore; hwang2022clinical). For instance, matthiesen2021clinician argued that, although XAI tools were perceived as potentially beneficial for transparency, they did not lead to substantial changes in clinical decision-making. Similarly, sivaraman2023ignore observed that XAI outputs were often regarded as supplementary evidence rather than as decisive information. Moreover, zhang2024rethinking noted that it may be more effective to decouple XAI from the final decision stage, positioning it instead as a supporting mechanism during model validation or clinician training. Furthermore, we also identify the sub-them (2.1) XAI not Always Actionable (bhattacharya2023directive; chari2023informing; ihongbe2024evaluating), where XAI often lacks clinical contextualization, making it insufficiently aligned with established diagnostic reasoning. Collectively, these findings suggest that while clinicians acknowledge the theoretical advantages of XAI, their reliance on established diagnostic reasoning and professional experience continues to dominate practical decision processes.
@@
Next, we identified Theme 3: Disparities Across Clinical Roles and Seniority Levelsgombolay2024effects; barda2020qualitative; cabitza2025fromoracular. A consistent observation across these studies was that clinicians’ experiences with CDSS varied according to their level of expertise. For instance, cabitza2025fromoracular reported that less experienced clinicians perceived the CDSS as cognitively demanding. Conversely,  gombolay2024effects found that more senior clinicians encountered greater difficulties integrating the system into their established workflows. Likewise, barda2020qualitative observed that clinical roles significantly shaped attitudes toward CDSS adoption, with preference given to solutions that minimized cognitive load and aligned with existing professional practices. These findings indicate that user acceptance and system usability in XAI-supported clinical environments are not uniform but mediated by clinicians’ roles, expertise, and cognitive expectations.
@@
Finally, we determined Theme 4: Frictions Between Stakeholders for DevelopmentBienefeld2023. Although only one study fits into this category, it is worth to discuss the implications of this theme. Here, Bienefeld2023 documented that frictions may arise between clinicians (as the CDSS users) and developers. Specifically, clinicians emphasized the importance of clinical plausibility and the alignment of model outputs with established medical reasoning, whereas developers prioritized model interpretability. Therefore, it may be logical to presume that early aligning stakeholders objectives is necessary to properly build usable CDSS without frictions. Contextually, we highlight that in only 12 studies (out of 31) early clinicians’ feedback was gathered before CDSS development barda2020qualitative; matthiesen2021clinician; Bienefeld2023; ellenrieder2023promoting; pumplun2023bringing; hwang2022clinical; zhang2024rethinking; bhattacharya2023directive; abraham2023integrating; gombolay2024effects; jing2025development. Early incorporation of clinicians’ perspectives can help ensure that system design and model outputs are aligned with real-world clinical expectations.
@@
1. Improvements in Risk Assessments brennan2019comparing he2024vms jung2025evaluating hur2025comparison kumarakulasinghe2020evaluating gu2020acase jing2025development rajashekar2024human famiglini2024evidence rainey2024operationalizing sabol2020explainable kayadibi2025ai singh2021evaluation chanda2024dermatologist pumplun2023bringing singla2023medical Documented Adoption Enhanced Diagnostics
 1.1 Enhanced Support for Novice Clinicians neves2021interpretable Better Novice Support
 1.2 Reduction of False Learning ellenrieder2023promoting Optimal Learning
 1.3 Enhancement of Clinical Planning abraham2023integrating Enhanced Planning
2. Preference for Established Clinical Practice kovalchuk2022three matthiesen2021clinician zhang2024rethinking sivaraman2023ignore hwang2022clinical Seek for Clinical Explanations Plausibility
 2.1 XAI not Always Actionable bhattacharya2023directive ihongbe2024evaluating chari2023informing Prefer Actionable Explanations
3. Disparities Across Clinical Roles / Seniority Levels gombolay2024effects barda2020qualitative cabitza2025fromoracular Implement Adaptive Explanations
4. Frictions Between Stakeholders for Development Bienefeld2023 Integrate Iterative Feedback Loops
@@ 4.2 A Socio-Technical Perspective for XAI-based CDSS
As previously discussed, Theme 1 can positively promote CDSS adoption in healthcare. However, Themes 2, 3, and 4 may generate frictions regarding CDSS usage. Under a theoretical perspective, such frictions may be attributed to socio-technical gaps between human requirements and the effective utility of technological solutions, because human activity is “flexible and nuanced”, while computational mechanisms are “rigid and brittle” ackerman2000intellectual. In other words, a socio-technical gap refers to the discrepancy between human needs, including expectations, and the technological capabilities designed to support them, highlighting the limitations of technological systems in fully accommodating the nuances of human interactions. This general theoretical construct can be extended to the purpose of XAI explanations in CDSS, as explanation capabilities might not meet rigorous human requirements from precise clinical knowledge (e.g., Theme 2), given that CDSS are dramatically at high stake, as output recommendations may have a significant impact on patient lives. Simply put, implementing XAI solutions in CDSS might not always bridge the socio-technical gap when compared to systems that do not adopt XAI, or systems that use basic reasoning in current clinical practice.
Therefore, based on the thematic analysis in Section 4.1, we suggest a reference protocol for effective CDSS development. This is an iterative process that requires exploring and connecting both the XAI and HCI fields, leveraging their theoretical constructs to eventually devise a practical methodology. We begin by identifying two theoretical dimensions: T1: determining stakeholders and their requirements for use case models (human requirement realism), and T2: developing XAI evaluation methods to reliably assess whether and to what extent these requirements are met in downstream use cases (context realism) liao2021human.
@@ 4.3 Stakeholder-Centric Protocol for CDSS Development
Core stakeholders include clinicians, developers, and healthcare organizations (hospitals, clinics, management). However, for real-world adoption, regulators, payers/insurers, and policymakers must also be considered shortliffe2018clinical. Stakeholder requirements are often divergent, as clinicians may emphasize improvements in risk assessment and XAI non-disruptive actionability (Themes 1,2), while managers may focus on efficiency, and regulators require compliance. This process inevitably necessitates to negotiate trade-offs vandepoel2013values. Although our review shows that empirical evaluations targeted clinicians, we argue that patients should also be considered as active stakeholders ancker2021invisible. Patients are directly affected by AI-informed decisions and increasingly encounter CDSS outputs in decision-making contexts. This absence highlights a non-negligible research gap: future work may design evaluation protocols that may incorporate patient perspectives, ensuring that explanations are not only useful by clinicians but also easily communicable to patients (when possible).
@@
Stakeholders would collaborate to define the clinical problem the CDSS will address to preventively predict any possible friction later (Theme 4). Shared problem framing ensures that system design aligns with actual decision-making needs rather than abstract performance benchmarks. Non-technical stakeholders must be informed of the CDSS’s underlying logic, including its machine learning models, XAI components, and evaluation frameworks. This stage establishes the evaluation criteria and metrics that guide subsequent development. To this end, Brankovic2025 formally proposed a reference set of 14 evaluation metrics for XAI in CDSS divided into clinical attributes, decision attributes, and model attributes that all stakeholders should be aware of.
@@
A progressive validation pipeline is recommended: initial testing with low-fidelity prototypes to explore the general structure, followed by retrospective studies in which clinicians may provide early feedback for more rapid iteration, and culminating in high-fidelity prototypes that closely resemble the final product in functionality. Prototypes should seamlessly integrate into existing clinical workflows (Theme 2) weir2021cognitive. Here, human-centered usability testing, such as think-aloud studies, can capture clinicians’ perceptions of system accuracy and explainability during interaction.
@@
  • Human-centered evaluation: trust, cognitive workload, workflow integration, and overall usability khairat2018reasons.
  • Ethical-legal considerations: accountability, transparency, and bias mitigation ghassemi2021false.
@@
Although the framework is field-agnostic, it requires adaptation to local contexts. For instance, radiology-oriented CDSS may emphasize imaging pipelines, whereas psychiatry may require interpretability in narrative or unstructured data. Hence, the framework is best understood as generalizable but context-sensitive, consistent with socio-technical perspectives on CDSS design greenes2014cds.

### v4

@@ 3.3 CDSS adopting XAI Methodologies in Deep Learning
Finally, across various clinical specialties, studies consistently highlight the importance of human-centered evaluation in determining the utility of XAI methods. In OphthalmologySingh et al. (2021) assessed multiple visual explanation techniques for a CNN-based model trained to diagnose retinal diseases using a counterfactual explanatory method, showing positive reception by radiologists. While in CardiologyNeves et al. (2021) similarly applied various explainability methods to a CDSS based on ECG time series data for predicting heart arrhythmias. In both domains, surveys identified attribution methods, specifically DeepTaylor and LIME, as particularly useful, especially for less experienced clinicians. In Critical CareZhang et al. (2024) developed SepsisLab, leveraging an attention module (explanatory part) feeding into an LSTM to highlight important variables for sepsis prediction. Here, semi-structured interviews revealed that while explanations were valued, shifting the AI focus away from final decision could further strengthen the user experience. Next, in DermatologyChanda et al. (2024) implemented GradCAM visualizations for CDSS assisting to diagnose skin melanoma. The system was evaluated with 116 clinicians across multiple phases. Although XAI did not significantly improve diagnostic accuracy, it notably increased clinicians’ confidence and trust. Finally, in Sleep MedicineHwang et al. (2022) developed a CNN-based CDSS to measure sleep staging, providing explanations through convolutional filters, saliency maps, and intermediate activations. User studies indicated that explanations effectively confirmed AI predictions but were less helpful in determining ambiguous cases, reflecting the clinicians’ desire to validate predictions while relying on their domain expertise.
@@
Of the 31 studies, we identified 14 studies leveraging deep learning for CDSS development. CAM-based methods (including GradCAM) were the most adopted XAI framework with 5 occurrences. However, we documented a scattered application of several frameworks ranging from integrated gradients to LIME, with no clear framework to dominate as compared to SHAP for CDSS evaluated in the previous section. As for clinicians’ perception, the majority proved to helpful with 11 studies showing positive reception Ellenrieder et al. (2023); Pumplun et al. (2023); Sabol et al. (2020); Kayadibi et al. (2025); Singh et al. (2021); Neves et al. (2021); Chanda et al. (2024); Singla et al. (2023); E. Ihongbe et al. (2024). The remaining papers documented neutral XAI perceptions with:  Cabitza et al. (2025) finding that less experienced practitioners found explanations cognitively demanding, Zhang et al. (2024) shifting XAI away from final decision was perceived more reliable, and  Hwang et al. (2022) finding that AI predictions were less helpful in revealing ambiguous cases, reflecting the clinicians’ preference to validate predictions while relying on their domain expertise.
@@ 4.1 Thematic Analysis of XAI-based CDSS
Next, we briefly outline the main evaluation themes that emerged, reporting a summary in Table 2. Firstly, positive perceptions were mostly linked to Theme 1: Improvements in Risk Assessments. Here, most of the works surveyed fall in this theme, e.g., Brennan et al. (2019); He et al. (2024); Hur et al. (2025); Kumar. et al. (2020); Jing et al. (2025); Famiglini et al. (2024); Kayadibi et al. (2025), among others, where XAI was seen as an augmentative assisting tool to clinicians to improve their diagnoses. In this theme, we identified sub-themes such as: (1.1) Support for Less Experienced Clinicians Neves et al. (2021), where junior clinicians perceived improvements in risk assessment, (1.2) Reduction of False Learning Ellenrieder et al. (2023), where clinicians were critically able to spot erroneous recommendations, and (1.3) Enhancement of Clinical Planning Abraham et al. (2023), implying that doctors could better schedule treatments regarding their patients.
@@
However, incorporating XAI may still pose challenges and limitations. Therefore, we discuss themes that emerged from studies expressing neutral perceptions of XAI. Firstly, several works align with Theme 2: Preference for Established Clinical Practice (Kovalchuk et al. (2022); Matthiesen et al. (2021); Zhang et al. (2024); Sivaraman et al. (2023); Hwang et al. (2022)). For instance, Matthiesen et al. (2021) argued that, although XAI tools were perceived as potentially beneficial for transparency, they did not lead to substantial changes in clinical decision-making. Similarly, Sivaraman et al. (2023) observed that XAI outputs were often regarded as supplementary evidence rather than as decisive information. Moreover, Zhang et al. (2024) noted that it may be more effective to decouple XAI from the final decision stage, positioning it instead as a supporting mechanism during model validation or clinician training. Furthermore, we also identify the sub-them (2.1) XAI not Always Actionable (Bhattacharya et al. (2023); Chari et al. (2023); E. Ihongbe et al. (2024)), where XAI often lacks clinical contextualization, making it insufficiently aligned with established diagnostic reasoning. Collectively, these findings suggest that while clinicians acknowledge the theoretical advantages of XAI, their reliance on established diagnostic reasoning and professional experience continues to dominate practical decision processes.
@@
Next, we identified Theme 3: Disparities Across Clinical Roles and Seniority LevelsGombolay et al. (2024); Barda et al. (2020); Cabitza et al. (2025). A consistent observation across these studies was that clinicians’ experiences with CDSS varied according to their level of expertise. For instance, Cabitza et al. (2025) reported that less experienced clinicians perceived the CDSS as cognitively demanding. Conversely,  Gombolay et al. (2024) found that more senior clinicians encountered greater difficulties integrating the system into their established workflows. Likewise, Barda et al. (2020) observed that clinical roles significantly shaped attitudes toward CDSS adoption, with preference given to solutions that minimized cognitive load and aligned with existing professional practices. These findings indicate that user acceptance and system usability in XAI-supported clinical environments are not uniform but mediated by clinicians’ roles, expertise, and cognitive expectations.
@@
Finally, we determined Theme 4: Frictions Between Stakeholders for DevelopmentBienefeld et al. (2023). Although only one study fits into this category, it is worth to discuss the implications of this theme. Here, Bienefeld et al. (2023) documented that frictions may arise between clinicians (as the CDSS users) and developers. Specifically, clinicians emphasized the importance of clinical plausibility and the alignment of model outputs with established medical reasoning, whereas developers prioritized model interpretability. Therefore, it may be logical to presume that early aligning stakeholders objectives is necessary to properly build usable CDSS without frictions. Contextually, we highlight that in only 12 studies (out of 31) early clinicians’ feedback was gathered before CDSS development Barda et al. (2020); Matthiesen et al. (2021); Bienefeld et al. (2023); Ellenrieder et al. (2023); Pumplun et al. (2023); Hwang et al. (2022); Zhang et al. (2024); Bhattacharya et al. (2023); Abraham et al. (2023); Gombolay et al. (2024); Jing et al. (2025). Early incorporation of clinicians’ perspectives can help ensure that system design and model outputs are aligned with real-world clinical expectations.
@@
1. Improvements in Risk Assessments Brennan et al. (2019) He et al. (2024) Jung et al. (2025) Hur et al. (2025) Kumar. et al. (2020) Gu et al. (2020) Jing et al. (2025) Rajashekar et al. (2024) Famiglini et al. (2024) Rainey et al. (2024) Sabol et al. (2020) Kayadibi et al. (2025) Singh et al. (2021) Chanda et al. (2024) Pumplun et al. (2023) Singla et al. (2023) Documented Adoption Enhanced Diagnostics
 1.1 Enhanced Support for Novice Clinicians Neves et al. (2021) Better Novice Support
 1.2 Reduction of False Learning Ellenrieder et al. (2023) Optimal Learning
 1.3 Enhancement of Clinical Planning Abraham et al. (2023) Enhanced Planning
2. Preference for Established Clinical Practice Kovalchuk et al. (2022) Matthiesen et al. (2021) Zhang et al. (2024) Sivaraman et al. (2023) Hwang et al. (2022) Seek for Clinical Explanations Plausibility
 2.1 XAI not Always Actionable Bhattacharya et al. (2023) E. Ihongbe et al. (2024) Chari et al. (2023) Prefer Actionable Explanations
3. Disparities Across Clinical Roles / Seniority Levels Gombolay et al. (2024) Barda et al. (2020) Cabitza et al. (2025) Implement Adaptive Explanations
4. Frictions Between Stakeholders for Development Bienefeld et al. (2023) Integrate Iterative Feedback Loops
@@ 4.2 A Socio-Technical Perspective for XAI-based CDSS
As previously discussed, Theme 1 can positively promote CDSS adoption in healthcare. However, Themes 2, 3, and 4 may generate frictions regarding CDSS usage. Under a theoretical perspective, such frictions may be attributed to socio-technical gaps between human requirements and the effective utility of technological solutions, because human activity is “flexible and nuanced”, while computational mechanisms are “rigid and brittle” Ackerman (2000). In other words, a socio-technical gap refers to the discrepancy between human needs, including expectations, and the technological capabilities designed to support them, highlighting the limitations of technological systems in fully accommodating the nuances of human interactions. This general theoretical construct can be extended to the purpose of XAI explanations in CDSS, as explanation capabilities might not meet rigorous human requirements from precise clinical knowledge (e.g., Theme 2), given that CDSS are dramatically at high stake, as output recommendations may have a significant impact on patient lives. Simply put, implementing XAI solutions in CDSS might not always bridge the socio-technical gap when compared to systems that do not adopt XAI, or systems that use basic reasoning in current clinical practice.
Therefore, based on the thematic analysis in Section 4.1, we suggest a reference protocol for effective CDSS development. This is an iterative process that requires exploring and connecting both the XAI and HCI fields, leveraging their theoretical constructs to eventually devise a practical methodology. We begin by identifying two theoretical dimensions: T1: determining stakeholders and their requirements for use case models (human requirement realism), and T2: developing XAI evaluation methods to reliably assess whether and to what extent these requirements are met in downstream use cases (context realism) Liao and Varshney (2021).
@@ 4.3 Stakeholder-Centric Protocol for CDSS Development
Core stakeholders include clinicians, developers, and healthcare organizations (hospitals, clinics, management). However, for real-world adoption, regulators, payers/insurers, and policymakers must also be considered Shortliffe and Sepúlveda (2018). Stakeholder requirements are often divergent, as clinicians may emphasize improvements in risk assessment and XAI non-disruptive actionability (Themes 1,2), while managers may focus on efficiency, and regulators require compliance. This process inevitably necessitates to negotiate trade-offs van de Poel (2013). Although our review shows that empirical evaluations targeted clinicians, we argue that patients should also be considered as active stakeholders Ancker et al. (2021). Patients are directly affected by AI-informed decisions and increasingly encounter CDSS outputs in decision-making contexts. This absence highlights a non-negligible research gap: future work may design evaluation protocols that may incorporate patient perspectives, ensuring that explanations are not only useful by clinicians but also easily communicable to patients (when possible).
@@
Stakeholders would collaborate to define the clinical problem the CDSS will address to preventively predict any possible friction later (Theme 4). Shared problem framing ensures that system design aligns with actual decision-making needs rather than abstract performance benchmarks. Non-technical stakeholders must be informed of the CDSS’s underlying logic, including its machine learning models, XAI components, and evaluation frameworks. This stage establishes the evaluation criteria and metrics that guide subsequent development. To this end, Brankovic and others (2025) formally proposed a reference set of 14 evaluation metrics for XAI in CDSS divided into clinical attributes, decision attributes, and model attributes that all stakeholders should be aware of.
@@
A progressive validation pipeline is recommended: initial testing with low-fidelity prototypes to explore the general structure, followed by retrospective studies in which clinicians may provide early feedback for more rapid iteration, and culminating in high-fidelity prototypes that closely resemble the final product in functionality. Prototypes should seamlessly integrate into existing clinical workflows (Theme 2) Weir et al. (2021). Here, human-centered usability testing, such as think-aloud studies, can capture clinicians’ perceptions of system accuracy and explainability during interaction.
@@
  • Human-centered evaluation: trust, cognitive workload, workflow integration, and overall usability Khairat et al. (2018).
  • Ethical-legal considerations: accountability, transparency, and bias mitigation Ghassemi et al. (2021).
@@
Although the framework is field-agnostic, it requires adaptation to local contexts. For instance, radiology-oriented CDSS may emphasize imaging pipelines, whereas psychiatry may require interpretability in narrative or unstructured data. Hence, the framework is best understood as generalizable but context-sensitive, consistent with socio-technical perspectives on CDSS design Greenes (2014).